# Fourier Learning Machines: Nonharmonic Fourier-Based Neural Networks for Scientific Machine Learning

### v1

@@ 5.2.2 Varying Initial Condition OCP
We solve the same OCP considered thus far, but now varying the initial condition $\boldsymbol{u_{0}}$ in a set $U_{0}$, defined as a disk on the 2–simplex, centered at $(0.20,0.20,0.60)$, with radius $0.15$. Three FLMs with with $N=27$ and three inputs $(t,u_{01},u_{02})$ (we recall that the third state variable and its initial value is dependent on the other two), are used to approximate the two state and one control variables. The three FLMs are trained on $250$ random points inside the disk, and then tested in $100$ newly generated random points on the disk with center at $(0.20,0.20,0.60)$. Given the cyclic symmetry of the 2–simplex and the underlying payoff matrix, two other disks, centered at $(0.20,0.60,0.20)$ and $(0.60,0.20,0.20)$, provided a sampling of 200 hundred extra points used as initial conditions for testing. This choice of testing points is a mere illustration on how one could leverage the inherent cyclic symmetry of this OCP in order to reduce the training space required to learn the approximated optimal control law. This is attained by informing the controller on what disk it is actuating on, allowing it to use the appropriate coordinate control learned at the disk used for training. Nonetheless, such a training choice does not undermine the results reported herein had we chosen to focus on a single disk in the 2–simplex.
The hyperparameters are tuned and a learning rate of $0.0005$ and beta values of 0.95 and 0.97, respectively, are selected for training. To quantify the approximation accuracy, we define the Mean Absolute Percentage Error (MAPE) on the objective function as follows:
@@
where $J^{FLM}_{j}$ and $J^{PMP}_{j}$ are the the objective function values obtained from the FLM–approximated solution and the PMP–based solution, respectively, for the $j^{th}$ test point. Over the test points on each of the three disks, The MAPEs are calculated to be $0.81\%$, $0.85\%$, and $0.72\%$ respectively, while the respective MAEs are calculated to be $2.54\times 10^{-3}$, $2.74\times 10^{-3}$ and $1.81\times 10^{-3}$. Fig. 9 demonstrates the percentage errors and the absolute errors (point–wise) in the test points in the three disks. These numerical results represent a substantial improvement from the ones initially reported in [33], where a finite Fourier series failed to approximate the optimal control law of this OCP in this case of varying initial condition.
Figure 9: (Left) Absolute percentage errors and (right) Absolute errors – in objective function value on test points within the circular disks on the simplex plane centered at $(0.2,0.2,0.6)$ and its cyclic permutations with a radius of $0.15$.
Our results show the FLM was capable of solving both PDEs and OCPs with the same level of accuracy of other traditional NN models. The OCP example showed that the FLM can learn to solve an entire family of OCPs, demonstrating its potential to tackle high–dimensional problems, thus our proposed nonharmonic Fourier series–inspired NN, the FLM, provides a robust and effective approach for a wide range of numerical computation for scientific problems.
@@ 6 Conclusions and Future Work
We have introduced the FLM, a neural architecture that learns optimal frequencies directly while systematically representing a multidimensional nonharmonic Fourier series, thereby bridging the gap between classical spectral methods and modern NN architectures. By jointly training frequencies, amplitudes, and phase shifts, FLMs overcome the restrictions of fixed harmonic Fourier series and adapts naturally to both periodic and nonperiodic problems. Its implementation as a NN allows it to leverage scientific computational and optimization tools, such as backpropagation, automatic differentiation and state–of–the–art solvers. Computational experiments on benchmark PDEs demonstrate that FLMs can match or outperform established architectures such as SIREN and vanilla feedforward NNs. Additionally, FLMs solve a family of OCPs with high accuracy, demonstrating it generalization capability and its potential for high–dimensional, nonlinear problems.
Future work will explore improved weight initialization strategies, potentially exploring how a customized approach (to each type of parameter – amplitude, frequency and bias), can improve convergence and training stability. Furthermore, the FLMs can be evaluated in the context of other application domains, such as signal processing, image reconstruction, spectral analysis etc. Lastly, the potential of FLMs to successfully solve different OCPs with more intricate constraints can still be further explored, for example, by considering OCPs with inequality constraints, bounded controls or customized constraints designed to enforce symmetries or any other known mathematical properties of either control or state space.
@@
$H_{\boldsymbol{n}}$ $=A_{1}^{(\boldsymbol{n})}\cos\left(n_{1}x_{1}+n_{2}x_{2}-\phi_{1}^{(\boldsymbol{n})}\right)+A_{2}^{(\boldsymbol{n})}\cos\left(n_{1}x_{1}-n_{2}x_{2}-\phi_{2}^{(\boldsymbol{n})}\right)$
$=A_{1}^{(\boldsymbol{n})}\left[\cos\phi_{1}^{(\boldsymbol{n})}\cos(n_{1}x_{1}+n_{2}x_{2})+\sin\phi_{1}^{(\boldsymbol{n})}\sin(n_{1}x_{1}+n_{2}x_{2})\right]$
$\quad+A_{2}^{(\boldsymbol{n})}\left[\cos\phi_{2}^{(\boldsymbol{n})}\cos(n_{1}x_{1}-n_{2}x_{2})+\sin\phi_{2}^{(\boldsymbol{n})}\sin(n_{1}x_{1}-n_{2}x_{2})\right]$
$=A_{1}^{(\boldsymbol{n})}\Big{[}\cos\phi_{1}^{(\boldsymbol{n})}\Bigl{(}\cos(n_{1}x_{1})\cos(n_{2}x_{2})-\sin(n_{1}x_{1})\sin(n_{2}x_{2})\Bigr{)}$
$\quad+\sin\phi_{1}^{(\boldsymbol{n})}\Bigl{(}\sin(n_{1}x_{1})\cos(n_{2}x_{2})+\cos(n_{1}x_{1})\sin(n_{2}x_{2})\Bigr{)}\Big{]}$
$\quad+A_{2}^{(\boldsymbol{n})}\Big{[}\cos\phi_{2}^{(\boldsymbol{n})}\Bigl{(}\cos(n_{1}x_{1})\cos(n_{2}x_{2})+\sin(n_{1}x_{1})\sin(n_{2}x_{2})\Bigr{)}$
$\quad+\sin\phi_{2}^{(\boldsymbol{n})}\Bigl{(}\sin(n_{1}x_{1})\cos(n_{2}x_{2})-\cos(n_{1}x_{1})\sin(n_{2}x_{2})\Bigr{)}\Big{]}$
$=\left(A_{1}^{(\boldsymbol{n})}\cos\phi_{1}^{(\boldsymbol{n})}+A_{2}^{(\boldsymbol{n})}\cos\phi_{2}^{(\boldsymbol{n})}\right)\cos(n_{1}x_{1})\cos(n_{2}x_{2})$
$\quad+\left(A_{1}^{(\boldsymbol{n})}\sin\phi_{1}^{(\boldsymbol{n})}-A_{2}^{(\boldsymbol{n})}\sin\phi_{2}^{(\boldsymbol{n})}\right)\cos(n_{1}x_{1})\sin(n_{2}x_{2})$
$\quad+\left(A_{1}^{(\boldsymbol{n})}\sin\phi_{1}^{(\boldsymbol{n})}+A_{2}^{(\boldsymbol{n})}\sin\phi_{2}^{(\boldsymbol{n})}\right)\sin(n_{1}x_{1})\cos(n_{2}x_{2})$
$\quad+\left(-A_{1}^{(\boldsymbol{n})}\cos\phi_{1}^{(\boldsymbol{n})}+A_{2}^{(\boldsymbol{n})}\cos\phi_{2}^{(\boldsymbol{n})}\right)\sin(n_{1}x_{1})\sin(n_{2}x_{2}).$
Comparing with (9) we get,
$a_{1}^{(\boldsymbol{n})}$ $=A^{(\boldsymbol{n})}_{1}\cos\phi_{1}^{(\boldsymbol{n})}+A^{(\boldsymbol{n})}_{2}\cos\phi_{2}^{(\boldsymbol{n})},\quad a_{2}^{(\boldsymbol{n})}=A^{\boldsymbol{(n)}}_{1}\sin\phi_{1}^{(\boldsymbol{n})}-A^{(\boldsymbol{n})}_{2}\sin\phi_{2}^{(\boldsymbol{n})},$
$a_{3}^{(\boldsymbol{n})}$ $=A^{(\boldsymbol{n})}_{1}\sin\phi_{1}^{(\boldsymbol{n})}+A^{(\boldsymbol{n})}_{2}\sin\phi_{2}^{(\boldsymbol{n})},\quad a_{4}^{(\boldsymbol{n})}=-A^{(\boldsymbol{n})}_{1}\cos\phi_{1}^{(\boldsymbol{n})}+A^{(\boldsymbol{n})}_{2}\cos\phi_{2}^{(\boldsymbol{n})}.$
Appendix B Numerical Results for Hyperparameter Selection

### v3

@@ 5.2.2 Varying Initial Condition OCP
We solve the same OCP considered thus far, but now varying the initial condition $\bm{u_{0}}$ in a set $U_{0}$, defined as a disk on the 2–simplex, centered at $(0.20,0.20,0.60)$, with radius $0.15$. Three FLMs with with $N=27$ and three inputs $(t,u_{01},u_{02})$ (we recall that the third state variable and its initial value is dependent on the other two), are used to approximate the two state and one control variables. The three FLMs are trained on $250$ random points inside the disk, and then tested in $100$ newly generated random points on the disk with center at $(0.20,0.20,0.60)$. Given the cyclic symmetry of the 2–simplex and the underlying payoff matrix, two other disks, centered at $(0.20,0.60,0.20)$ and $(0.60,0.20,0.20)$, provided a sampling of $200$ extra points used as initial conditions for testing. This choice of testing points is a mere illustration on how one could leverage the inherent cyclic symmetry of this OCP in order to reduce the training space required to learn the approximated optimal control law. This is attained by informing the controller on what disk it is actuating on, allowing it to use the appropriate coordinate control learned at the disk used for training. Nonetheless, such a training choice does not undermine the results reported herein had we chosen to focus on a single disk in the 2–simplex.
The hyperparameters are tuned and a learning rate of $5\times 10^{-4}$ and beta values of 0.95 and 0.97, respectively, are selected for training. To quantify the approximation accuracy, we define the Mean Absolute Percentage Error (MAPE) on the objective function as follows:
@@
where $J^{FLM}_{j}$ and $J^{PMP}_{j}$ are the objective function values obtained from the FLM–approximated solution and the PMP–based solution, respectively, for the $j^{th}$ test point. Over the test points on each of the three disks, the MAPEs are calculated to be $0.81\%$, $0.85\%$, and $0.72\%$ respectively, while the respective MAEs are calculated to be $2.54\times 10^{-3}$, $2.74\times 10^{-3}$ and $1.81\times 10^{-3}$. Fig. 8 demonstrates the percentage errors and the absolute errors (point–wise) in the test points in the three disks. These numerical results represent a substantial improvement from the ones initially reported in Nicolosi (2023), where a finite Fourier series failed to approximate the optimal control law of this OCP in this case of varying initial condition.
Figure 8: (Left) Absolute percentage errors and (right) absolute errors in the objective function value on test points within the circular disks on the simplex plane centered at $(0.2,0.2,0.6)$ and its cyclic permutations with a radius of $0.15$.
5.2.3 Varying Initial Condition 5–dimensional OCP
To demonstrate the scalability of FLMs, we solve the 5–strategy extension of RPS, known as RPSSL (Rock–Paper–Scissors–Spock–Lizard). We extend the OCP in Equation 5.2 by replacing the matrices $L_{3}$ and $M_{3}$, respectively, in Equation 17 with
$\bm{L}_{5}=\begin{pmatrix}0&-1&1&-1&1\\ 1&0&-1&1&-1\\ -1&1&0&-1&1\\ 1&-1&1&0&-1\\ -1&1&-1&1&0\\ \end{pmatrix},\quad\bm{M}_{5}=\begin{pmatrix}0&0&1&0&1\\ 1&0&0&1&0\\ 0&1&0&0&1\\ 1&0&1&0&0\\ 0&1&0&1&0\end{pmatrix}$.
For this experiment, $4$ FLMs with $5$ inputs and $N=16$ sub–networks were used to approximate $4$ states and an FLM with $5$ inputs and $N=32$ sub–networks was employed to approximate the optimal control. In this case, the equilibrium point is $\bm{u}_{eq}=(0.2,0.2,0.2,0.2,0.2)$. All of the FLMs were trained and tested following the same method discussed before with a learning rate of $10^{-4}$ and beta values of $0.95$ and $0.97.$ Training the models with initial conditions (uniformly) sampled from a hyperdisk in the 44-simplex $\Delta_{4}$ of radius 0.0450.045111This choice of radius length is such that the density of sampled initial conditions is the same as in the RPS case. centered at $(0.11,0.11,0.11,0.11,0.56),$ we have achieved an MAPE of $0.67\%$ with MAE of $2.10\times 10^{-3}.$
Our results show that FLMs are capable of numerically solving PDEs with the same level of accuracy of other traditional NN models, often outperforming them in terms of error metrics considered. The OCP example also illustrated how FLMs can approximate the state trajectory and the optimal control of a high–dimensional family of OCPs.
@@ 6 Conclusions and Future Work
We have introduced the FLM, a neural architecture that learns optimal frequencies directly while systematically representing a multidimensional nonharmonic Fourier series, thereby bridging the gap between classical spectral methods and modern NN architectures. By jointly training frequencies, amplitudes, and phase shifts, FLMs overcome the restrictions of fixed harmonic Fourier series and adapts naturally to both periodic and nonperiodic problems. Its implementation as a NN allows it to leverage scientific computational and optimization tools, such as backpropagation, automatic differentiation and state–of–the–art solvers. Computational experiments on benchmark PDEs demonstrate that FLMs can match or outperform established architectures such as SIREN and vanilla feedforward NNs. Additionally, FLMs solve a family of OCPs with high accuracy, demonstrating its generalization capability and its potential for high–dimensional, nonlinear problems.
Future work will investigate an improved initialization scheme for the FLM weights (frequencies and amplitudes), potentially important for when the target function is dominated by high–frequency components. Additionally, one can consider applying FLMs to other domains, such as image and audio processing, medical signal processing, and time series analysis (a brief time series imputation study is presented in Appendix B). Finally, deeper theoretical relations, in particular with SSGP (Lázaro-Gredilla et al. (2010)) can be further investigated, potentially strengthening and better understanding the underlying principles of the spectral adaptability of both FLMs and SSGP models.
Appendix
@@
$H_{\bm{n}}$ $=A_{1}^{(\bm{n})}\cos\left(n_{1}x_{1}+n_{2}x_{2}-\phi_{1}^{(\bm{n})}\right)+A_{2}^{(\bm{n})}\cos\left(n_{1}x_{1}-n_{2}x_{2}-\phi_{2}^{(\bm{n})}\right)$
$=A_{1}^{(\bm{n})}\left[\cos\phi_{1}^{(\bm{n})}\cos(n_{1}x_{1}+n_{2}x_{2})+\sin\phi_{1}^{(\bm{n})}\sin(n_{1}x_{1}+n_{2}x_{2})\right]$
$\quad+A_{2}^{(\bm{n})}\left[\cos\phi_{2}^{(\bm{n})}\cos(n_{1}x_{1}-n_{2}x_{2})+\sin\phi_{2}^{(\bm{n})}\sin(n_{1}x_{1}-n_{2}x_{2})\right]$
$=A_{1}^{(\bm{n})}\Big[\cos\phi_{1}^{(\bm{n})}\Bigl(\cos(n_{1}x_{1})\cos(n_{2}x_{2})-\sin(n_{1}x_{1})\sin(n_{2}x_{2})\Bigr)$
$\quad+\sin\phi_{1}^{(\bm{n})}\Bigl(\sin(n_{1}x_{1})\cos(n_{2}x_{2})+\cos(n_{1}x_{1})\sin(n_{2}x_{2})\Bigr)\Big]$
$\quad+A_{2}^{(\bm{n})}\Big[\cos\phi_{2}^{(\bm{n})}\Bigl(\cos(n_{1}x_{1})\cos(n_{2}x_{2})+\sin(n_{1}x_{1})\sin(n_{2}x_{2})\Bigr)$
$\quad+\sin\phi_{2}^{(\bm{n})}\Bigl(\sin(n_{1}x_{1})\cos(n_{2}x_{2})-\cos(n_{1}x_{1})\sin(n_{2}x_{2})\Bigr)\Big]$
$=\left(A_{1}^{(\bm{n})}\cos\phi_{1}^{(\bm{n})}+A_{2}^{(\bm{n})}\cos\phi_{2}^{(\bm{n})}\right)\cos(n_{1}x_{1})\cos(n_{2}x_{2})$
$\quad+\left(A_{1}^{(\bm{n})}\sin\phi_{1}^{(\bm{n})}-A_{2}^{(\bm{n})}\sin\phi_{2}^{(\bm{n})}\right)\cos(n_{1}x_{1})\sin(n_{2}x_{2})$
$\quad+\left(A_{1}^{(\bm{n})}\sin\phi_{1}^{(\bm{n})}+A_{2}^{(\bm{n})}\sin\phi_{2}^{(\bm{n})}\right)\sin(n_{1}x_{1})\cos(n_{2}x_{2})$
$\quad+\left(-A_{1}^{(\bm{n})}\cos\phi_{1}^{(\bm{n})}+A_{2}^{(\bm{n})}\cos\phi_{2}^{(\bm{n})}\right)\sin(n_{1}x_{1})\sin(n_{2}x_{2}).$
Comparing with Equation 9 we get,
$a_{1}^{(\bm{n})}$ $=A^{(\bm{n})}_{1}\cos\phi_{1}^{(\bm{n})}+A^{(\bm{n})}_{2}\cos\phi_{2}^{(\bm{n})},\quad a_{2}^{(\bm{n})}=A^{\bm{(n)}}_{1}\sin\phi_{1}^{(\bm{n})}-A^{(\bm{n})}_{2}\sin\phi_{2}^{(\bm{n})},$
$a_{3}^{(\bm{n})}$ $=A^{(\bm{n})}_{1}\sin\phi_{1}^{(\bm{n})}+A^{(\bm{n})}_{2}\sin\phi_{2}^{(\bm{n})},\quad a_{4}^{(\bm{n})}=-A^{(\bm{n})}_{1}\cos\phi_{1}^{(\bm{n})}+A^{(\bm{n})}_{2}\cos\phi_{2}^{(\bm{n})}.$
Appendix B FLM For Time Series Imputation
To demonstrate the potential of FLMs for time series imputation tasks, we present in this Appendix some numerical results derived from a brief set of computational experiments using data from the Electricity Transformer Temperature (ETT) dataset Zhou et al. (2021). This dataset is widely used in time series forecasting and imputation benchmarks due to its strong quasi–periodic patterns and significant fluctuations, making it a good testbed candidate to evaluate the frequency–learning capabilities of our model. We compare the performance of FLMs to the models presented in Du et al. (2024), also evaluated for the ETT data set.
B.1 Dataset and Experimental Setup
We utilize the ETTh1 (Electricity Transformer Temperature – 1 hour) subset, which contains hourly records of oil temperature (OT) and other load features from electricity transformers over a period of two years. For this experiment, we focus on the univariate imputation of the OT channel only, which exhibits both daily and yearly seasonality alongside high–frequency oscillations.
We replicate the data preprocessing and experimental protocol present in Du et al. (2024):
  • Preprocessing: The data is normalized using a standard scaler (zero mean, unit variance). The time index $t$ is mapped to the domain $\pi,\pi]$.
  • Splitting: We employ a chronological split of 60% for training, 20% for validation, and 20% for testing.
  • Masking: To simulate missing values, we apply random point-wise masking at two different missingness rates: 10% and 50%.
  • Training: The model is trained to minimize the MSE on the observed (non-masked) data points of the training set. After hyperparameter tuning, we utilize the Adam optimizer with a learning rate of $1\times 10^{-4}$ and a hidden layer size of $N=8$. The training runs for 10,000 epochs with early stopping based on validation loss.
B.2 Results
We report the FLM approximation error on the held-out masked values of the test set and the performance is evaluated using MAE. Table 3 summarizes the results averaged over 10 independent runs with different random seeds. For comparative baselines, we refer to the comprehensive results established in Du et al. (2024), noting that the FLM achieved a competitive performance (although not superior) with a significantly fewer number of parameters and much simpler architecture than most of the models considered in Du et al. (2024) (mostly Deep NN models). In particular, referring to the performance of models in Du et al. (2024) for the dataset considered, FLMs would rank $26$ and $24$ out of $29$ for the cases of $10\%$ and $50\%$ of point–wise missing data, respectively. We emphasize that such performance was attained with a single–input FLM only. These preliminary results indicate that FLMs have great potential to be explored in this domain, and perhaps can be further studied or adapted for the specific task of time series imputation.
Table 3: FLM imputation results on ETTh1 (Mean $\pm$ SD over 10 seeds) against some of the performance of state-of-the-art models considered in Du et al. (2024) (** refers to the top performing and * to the lowest ranked one).
MAE
Model Mean (Standard Deviation)
10% Missing Rate
FLM 0.57 (0.15)
MRNN* (Yoon et al. (2018)) 0.79 (0.02)
Mean (Du et al. (2024)) 0.74 (–)
Median (Du et al. (2024)) 0.71 (–)
SAITS** (Du et al. (2023)) 0.14 (0.01)
50% Missing Rate
FLM 0.66 (0.18)
AutoFormer* (Wu et al. (2021)) 0.98 (0.01)
Mean (Du et al. (2024)) 0.74 (–)
Median (Du et al. (2024)) 0.71 (–)
SAITS** 0.22 (0.01)
Appendix C Numerical Results for Hyperparameter Selection
When numerically solving each PDE, we trained FLMs and other NN models for 18–27 different hyperparameter configurations, with a maximum of 10,000 epochs and a training loss tolerance of $10^{-4}$. All error metrics (MSE, MAE, Max Error) and the number of epochs required to reach the specified training loss tolerance were averaged over 10 runs with different random seeds. The performances of all tested models applied to each PDE are reported in Tables 526. Table 4 shows the number of trainable parameters in each model for different number of hidden units. The effect of the number of sub–networks in approximating the solutions of the PDEs is shown in figure Fig. 9.
Table 4: Number of trainable parameters for different models. For FLM, NN denotes the number of sub–networks. For all other models, $N_{\text{hidden}}$ denotes the number of neurons in each hidden layer. For each row, parameter counts are reported in the same order as the listed NN or $N_{\text{hidden}}$ values.
Model Hidden Units Formula Trainable Parameters
FLM $\qquad N\qquad=\{04,\,16,\,25,\,49,\,64\}$ $6N$ $\{24,\,96,\,150,\,294,\,384\}$
SIREN–1 $N_{\text{hidden}}=\{06,\,24,\,37,\,73,\,93\}$ $4N_{\text{hidden}}+1$ $\{25,\,97,\,149,\,293,\,373\}$
SIREN–2 $N_{\text{hidden}}=\{03,\,08,\,10,\,15,\,17\}$ $N_{\text{hidden}}^{2}+5N_{\text{hidden}}+1$ $\{25,\,105,\,151,\,301,\,375\}$
SIREN–3, V-ReLU, V-LReLU, V-Tanh $N_{\text{hidden}}=\{02,\,06,\,07,\,11,\,12\}$ $2N_{\text{hidden}}^{2}+6N_{\text{hidden}}+1$ $\{21,\,109,\,141,\,309,\,361\}$
Figure 9: Effect on approximation MSE for the PDEs considered in Section 5 (shown in vertical axis in logarithmic scale) as NN increases. All FLMs were trained using default learning rate and beta values of ADAM optimizer in PyTorch with maximum epochs of $10,000$ and loss tolerance of $10^{-4}.$
With the best performing hyperparameter configuration (highlighted in each table below) selected by following the aforementioned procedure, we train the NN models for $30,000$ new additional epochs with a training loss tolerance of $10^{-8}$, and report the error metrics for each PDE in their respective Tables 2(a)2(d).